OpenStudy (anonymous):

4(x+1)-(3x+5)=1 Somehow, I keep getting x=-8. My order of operations is breaking down. Can anyone explain what I'm doing wrong? 4x+4-3x+5=1 x+9=1 x=-8

OpenStudy (anonymous):

3x + 6 = 4x + 4 When you solve this, you get x = 2 as 3x - 4x = -2, -x = -2, x = 2

OpenStudy (anonymous):

its -5 in the second step not +5

OpenStudy (ranga):

@j.thain as rizwan pointed out, the mistake is in your first step: the +5 should be -5. 4x + 4 - 3x - 5 = 1 x - 1 = 1 x = 2